40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 31 10:16:02 2012, janosch, Steve, Update, , ETMX, scattering preps IMG_1609.JPGIMG_1608b.jpg
    Reply  Fri Sep 7 09:37:20 2012, Steve, Update, Cameras, baffle plate for SOS  IMG_1624.JPGIMG_1618.JPGIMG_1616.JPG
       Reply  Wed Sep 12 17:02:00 2012, Steve, Update, Cameras, baffle plate hole getting larger IMG_1631.JPGIMG_1628.JPG
          Reply  Thu Nov 21 09:56:15 2013, Steve, Update, SUS, green glass beam dumps HEXdump.jpgVdump.jpgDiamond1_2inchTraps.jpg
             Reply  Thu Nov 21 10:24:50 2013, Koji, Update, SUS, beam dumps 
Message ID: 7326     Entry time: Fri Aug 31 10:16:02 2012     Reply to this: 7358
Author: janosch, Steve 
Type: Update 
Category:  
Subject: ETMX, scattering preps 

The alignment of the pick-off mirror near ETMX is done. Everything turned out to be easy once we realized that there is no sense getting the alignment laser (going through viewport to pick-off to ITMX) back to ETMX. It is only necessary to hit ITMX somehow, since this makes sure that there is one scattered beam that will make it from ITMX to pick-off through viewport.

After the auxiliary optic (that we never used in the end) was removed again, we levelled the optical table.

So in the current setup, we can have small-angle scattering measurements on ITMX and large-angle scattering measurements on ETMX.

Attachment 1: IMG_1609.JPG  2.055 MB  Uploaded Fri Aug 31 11:36:06 2012  | Hide | Hide all | Show all
IMG_1609.JPG
Attachment 2: IMG_1608b.jpg  2.073 MB  Uploaded Tue Sep 4 18:27:21 2012  | Show | Hide all | Show all
ELOG V3.1.3-